OpenStudy (anonymous):

prove the following identities csc(theta) - sin(theta) = cos(theta) * cot (theta)

OpenStudy (anonymous):

1/sin - sin =(1-sin^2) /sin = cos^2 / sin = (cos/sin)cos=cot cos